Session: Feb 2012 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (Uwe Bach at Thu Mar 01 09:27:56 2012) Failure:
Not yet done because of unfavourable weather. Next try 13./14.3. together with CL12C1.

Hartebeesthoek (Jonathan Quick at Thu Mar 01 12:51:53 2012) Failure:
Not done - SEFD too high.

Onsala (Michael Lindqvist at Mon Mar 05 10:59:17 2012) Success:
We don't have a noise diode to calibrate.

Metsahovi (E. Rastorgueva at Mon Mar 05 11:56:26 2012) Success:
No known problem

Effelsberg (Uwe Bach at Thu Mar 15 10:20:11 2012) Success:
Now done at clear sky conditions!

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:15:13 2012) Success:
Not done, Rx not sensitive enough...